OpenStudy (ashleyvaldes501):

Which of the following is a solution of 3x2 = 7x − 3?

jhonyy9 (jhonyy9):

following - where are ?

OpenStudy (ashleyvaldes501):

negative 7 plus or minus the square root of 13 divided by 6 7 plus or minus the square root of 85 divided by 6 negative 7 plus or minus the square root of 85 divided by 6 7 plus or minus the square root of 13 divided by 6

jhonyy9 (jhonyy9):

do you know how you can solving quadratic equation ?

OpenStudy (ashleyvaldes501):

whats the equation?

jhonyy9 (jhonyy9):

3x^2 = 7x -3 what you ve wrote above

jhonyy9 (jhonyy9):

take the 7x-3 on the left side - this is the first step

OpenStudy (ashleyvaldes501):

u equal it to zero?

OpenStudy (ashleyvaldes501):

ok after are u to factor?

jhonyy9 (jhonyy9):

after take the 7x-3 on the left side yes will make it equal zero

jhonyy9 (jhonyy9):

do you can writing me here please how you ve get it ?

jhonyy9 (jhonyy9):

and after this you need to calculi the discriminant - do you know the formula of discriminant ?

OpenStudy (ashleyvaldes501):

b^2-4ac

jhonyy9 (jhonyy9):

yes right but what is your quadratic firstly ?

OpenStudy (ashleyvaldes501):

oh ok I'm getting it let me solve it

jhonyy9 (jhonyy9):

3x^2 = 7x -3 subtract from both sides 7x and add to the both sides 3

OpenStudy (ashleyvaldes501):

3x^2-7x+3, then i just put it into quad form? a=3,b=-7,c=3??

jhonyy9 (jhonyy9):

exactly - very nice job - but 3x^2 -7x +3 make it equal zero - ok. ?

OpenStudy (ashleyvaldes501):

yes I got it now, so its the third one, thank you so much!

jhonyy9 (jhonyy9):

negative 7 plus or minus the square root of 85 divided by 6 do you think on this ?

jhonyy9 (jhonyy9):

3x^2-7x+3, then i just put it into quad form? a=3,b=-7,c=3?? so in this case just you need calculi the discriminant using these a,b and c values and this above wrote formula of discriminant

OpenStudy (ashleyvaldes501):

never mind that answer was wrong, not sure where i made my mistake i just put it onto quad form and solved

jhonyy9 (jhonyy9):

yes i saw it just for tis i wan ,like solve it here and i can checking your work - ok ?

jhonyy9 (jhonyy9):

Ashley - how many is the discriminant - this is the next step

jhonyy9 (jhonyy9):

3x^2-7x+3, then i just put it into quad form? a=3,b=-7,c=3?? you wrote this above

OpenStudy (ashleyvaldes501):

3x^2-7x+3 quad form: x=-(-7)+/- sqrt of -7^2 -4(3)(3) so its 7+/-sqrt of 49-36 7+/-sqrt of 13

OpenStudy (ashleyvaldes501):

so its the last one but idk where they got the 6 from

jhonyy9 (jhonyy9):

Ashley - sorry but i ve wrote above that you need make it equal zero for you get an equation - 3x^2 -7x +3 = 0 D = b^2 -4ac = ?

OpenStudy (ashleyvaldes501):

-7^2-4(3)(3)

OpenStudy (ashleyvaldes501):

49-36

jhonyy9 (jhonyy9):

yes for discriminant is right so how many will be ?

OpenStudy (ashleyvaldes501):

13

jhonyy9 (jhonyy9):

correct now you know that the discriminant is equal 13 and you need to know formula of roots x_1,2 = ?

OpenStudy (ashleyvaldes501):

idk the formula

jhonyy9 (jhonyy9):

you know it sure bc. you ve wrote it above but you forgot something from there - check it please

jhonyy9 (jhonyy9):

there you ve wrote x = -(-7) +/- ...

jhonyy9 (jhonyy9):

so bc. there are +/- this mean that is formula of two roots x_1,2 = ...

OpenStudy (ashleyvaldes501):

oh ok i wasn't sure what you were asking, so 49-sqrt13 49+sqrt13

jhonyy9 (jhonyy9):

from where is this 49 there ?

OpenStudy (ashleyvaldes501):

oh I'm sorry i was squaring it.. I'm all over the place sorry.. then 7-sqrt13 7+sqrt13

jhonyy9 (jhonyy9):

Ashley - just easy . please relax one moment - you know these all

jhonyy9 (jhonyy9):

ok. looke please there i said that this what you ve wrote there not is the complete answer bc you forgot ,missed from there the last part of this formula

jhonyy9 (jhonyy9):

please you dont remember x_1,2 = (-b +/- sqrt(D)) .?... what is there again ?

jhonyy9 (jhonyy9):

no please delete it

jhonyy9 (jhonyy9):

so you know that there is x_1,2 = (-b +/- sqrt(D)) .?.. and from the end of this formula you ve forgot something

OpenStudy (ashleyvaldes501):

omg /2a

jhonyy9 (jhonyy9):

than i said that you need divide it by something

jhonyy9 (jhonyy9):

yes sure - do you see that you know these ALL

OpenStudy (ashleyvaldes501):

haha wow i can't believe this is what i have been forgetting this entire time.. and that equals 6.. wow seriously thank you so much, i appreciate it a lot!

jhonyy9 (jhonyy9):

Ashley i said to you that you know these ALL just you need relax one moment and hope these helped to you was my pleasure and anytime just tag me than you like help you - so than i m here online good luck bye bye

jhonyy9 (jhonyy9):

or do you have again questions ?

OpenStudy (ashleyvaldes501):

no haha thats it for now if i have any other questions ill be sure to tag you! :)

jhonyy9 (jhonyy9):

ok will be here and help you - hope i will can - bye bye